OpenStudy (anonymous):

Carlos wants to deposit $900 into savings accounts at three different banks: Bank of Chance, Merchant Bank, and Utopian Financing. He will deposit two times as much into Merchant Bank as Bank of Chance because they offer a higher interest rate. He also expects the Utopian Financing deposit to be only 20% of the total of the other two deposits. How much will Carlos deposit into the Utopian Financing savings account? $150 $180 $250 $500

OpenStudy (anonymous):

Bank of Chance: $900 Merchant Bank: $1800

OpenStudy (anonymous):

So 900 + 1800 = 2700

OpenStudy (anonymous):

.2 * 2700 = 540

OpenStudy (anonymous):

hmm I did something wrong :S

OpenStudy (zpupster):

x in Bank of Chance 2x in Merchant Bank these equal 3x and .2 or 1/5 in of the amount in those 2 banks so 3x + 3x/5 = 900 when you get x add x and 2x together and multiply by .2 or divide by 5